OpenStudy (anonymous):

Find an equation for the parabola and put it in the form: y=a(x-h)^2+k vertex= 3,1 point on graph= 4,2

OpenStudy (anonymous):

2 = a(4 - 3)^2 + 1, solve for a

OpenStudy (anonymous):

what about this one? vertex=1,-1 point on graph= 0,-2

OpenStudy (anonymous):

@xxferrocixx please put this in the equation form.